# Joint Graph Rewiring and Feature Denoising via Spectral Resonance

### v2

@@ 3 Experiments
Table 1: Comparison of state-of-the-art preprocessing rewiring approaches. Note that we refer to the complexity per iteration. $N$ denotes the number of nodes, $m$ the number of edges and $d_{\text{max}}$ is the maximum node degree.
@@
Method Add edge Remove edge Use Features Heterophilic? Complexity
DIGL (Gasteiger et al., 2019) $\mathcal{O}(N)$
FoSR (Karhadkar et al., 2023) $\mathcal{O}(N^{2})$
BORF (Nguyen et al., 2023) $\mathcal{O}(md_{\text{max}}^{3})$
JDR (Ours) $\mathcal{O}(N)$
@@
We extensively evaluate JDR on both synthetic data generated from the cSBM and real-world benchmark datasets. We follow experimental setting from Chien et al. (2021) and evaluate JDR for semi-supervised node classification with different downstream GNNs. We also adopt their data splits, namely the sparse splitting $2.5\%/2.5\%/95\%$ for training, validation and testing, respectively, or the dense splitting $60\%/20\%/20\%$. For the general experiments, we perform $100$ runs with different random splits. For the scalability experiments, we use the experimental settings of the respective works (Lim et al., 2021; Platonov et al., 2023). We report the average accuracy and the $95\%$-confidence interval calculated via bootstrapping with 1000 samples. All experiments are reproducible using the code provided.
@@
Baselines. Following recent works on rewiring, we use graph convolution network (GCN) (Kipf and Welling, 2017) as our downstream GNN. To obtain a more comprehensive picture, we additionally evaluate the performance on the more recent generalized PageRank graph neural network (GPRGNN) (Chien et al., 2021). We compare our algorithm with the state-of-the-art rewiring methods first-order spectral rewiring (FoSR) (Karhadkar et al., 2023), batch Ollivier-Ricci flow (BORF) (Nguyen et al., 2023) and diffusion improves graph learning (DIGL) (Gasteiger et al., 2019). FoSR approximates which edges should be added to maximize the spectral gap to reduce oversquashing. BORF adds edges in regions of negative curvature in the graph, which indicate bottlenecks that can lead to an oversquashing of the messages passed along these edges. A positive curvature indicates that there are so many edges in this area that messages could be oversmoothed, which is why edges are removed here. We compare computational and implementation aspects of JDR and baselines in Table 1. On the cSBM, we compare to an optimal baseline, namely the approximate message passing-belief propagation (AMP-BP) algorithm (Duranthon and Zdeborová, 2023). AMP-BP is asymptotically optimal (in the large dimension limit) for unsupervised or semi-supervised community detection in the cSBM. It relies on knowing the distribution of the cSBM and is thus not applicable to real-world graphs with unknown characteristics and complex features.
@@
Hyperparameters. Unless stated otherwise, we use the hyperparameters from Chien et al. (2021) for the GNNs and optimize the hyperparameters of JDR using a mixture of grid and random search on the validation set. We use the top-$64$ values of $\tilde{\bm{A}}$ to enforce sparsity and interpolation to update the features. For DIGL, FoSR and BORF, we tune their hyperparameters using a grid search, closely following the given parameter range from the original papers. For all hyperparameter searches we use GCN and GPRGNN as the downstream models on $10$ runs with different random splits. A detailed list of all hyperparameters can be found in Appendix A.7 or in the code repository.
@@ 3.2 Results on Real-World Data
We evaluate JDR on five common homophilic benchmarks datasets, namely the citation graphs Cora, CiteSeer, PubMed (Sen et al., 2008) and the Amazon co-purchase graphs Computers and Photo (McAuley et al., 2015). For heterophilic datasets, we rely on the Wikipedia graphs Chameleon and Squirrel (Rozemberczki et al., 2021), the WebKB datasets Texas and Cornell used in Pei et al. (2020) and the actor co-occurence network Actor (Tang et al., 2009). To show the scalability of JDR on larger heterophilic datasets, we further report the results for the Yandex Q user network Questions (Platonov et al., 2023) and the social networks Penn94 and Twitch-Gamers (Lim et al., 2021). Further details about all datasets are in Appendix A.4. Following Chien et al. (2021), we evaluate the homophilic datasets in the sparse splitting, staying close to the original setting of Kipf and Welling (2017) and the heterophilic datasets in dense splitting (Pei et al., 2020). The remaining larger graphs are evaluated using their original splits. For further results and splits, see Appendix A.5.
@@
Heterophilic Datasets. The results in Table 3 show that GCN +JDR can catch up significantly compared to GPRGNN, but GPRGNN +JDR generally performs better. This is in line with the findings for the heterophilic cSBM. DIGL performs well on Actor, Texas and Cornell despite its inherent homophily assumption. The reason for this is the chosen smoothing kernel, which results in a graph that is evenly connected everywhere with small weights. GCN then largely ignores the graph and thus performs very similarly to an MLP, which performs already quite well on these datasets (Chien et al., 2021). However, this fails for GPRGNN, which can make better use of the weak, complex graph structures. FoSR and BORF also improve performance here in most cases, but they are outperformed by JDR in all cases, often by a large margin. The out-of-memory error on Squirrel for BORF results from its computational complexity of $\mathcal{O}(md_{\text{max}}^{3})$, because the dataset has a large number of edges $m$ and a high maximum node degree $d_{\text{max}}$.
@@ 4 Relation to Prior Work
Graph Rewiring. Recent work show that even when a graph correctly encodes interactions it may not be an effective computational graph for a GNN due to conditions such as oversquashing (Alon and Yahav, 2021; Di Giovanni et al., 2023) and oversmoothing (Chen et al., 2020a). Recently many methods have been proposed to address this, notably graph rewiring methods. They can be divided into preprocessing and end-to-end methods. Preprocessing methods rewire the graph using geometric and spectral properties, including curvature (Topping et al., 2022; Nguyen et al., 2023; Fesser and Weber, 2024; Bober et al., 2024), expansion (Deac et al., 2022; Banerjee et al., 2022), effective resistance (Black et al., 2023; Shen et al., 2024), and spectral gap (Karhadkar et al., 2023). Conceptually related is diffusion-based rewiring (Gasteiger et al., 2019) that smooths the graph with a diffusion kernel. This can be interpreted as graph denoising, but is only suitable for homophilic graphs. Our approach is related to rewiring but with several key differences (see Table 1). Our rewiring strategy aims to denoise the graph (rather than control some geometric property) with the goal to improve downstream performance, while the classical rewiring literature focuses on optimizing the graph for message passing computations.
@@
Early end-to-end methods randomly drop edges during training to reduce oversmoothing (Rong et al., 2020). Subsequent work (Gutteridge et al., 2023; Qian et al., 2024) incorporates latent features to dynamically rewire the graph. Ji et al. (2023) use the estimated labels of a GNN to rewire the graph during training of the same GNN. Giraldo et al. (2023) use curvature information for dynamic rewiring. Graph Transformers (Dwivedi and Bresson, 2021; Rampasek et al., 2022) aim to overcome oversquashing in GNNs via global attention. In order to handle large graphs, these works still need to revert to sparse, non-global attention (Gabrielsson et al., 2023; Shirzad et al., 2023). Unlike preprocessing methods, end-to-end methods cannot output an improved graph which restricts their interpretability and reusability.
Graph Denoising. There is extensive literature on denoising signals on graphs using graph filters (Chen et al., 2014; Ma et al., 2021b; Liu et al., 2022). However, we are interested in modifying the structure of the graph itself (rewiring), in a way that can benefit any downstream algorithm.
@@
Graph Structure Learning. The aim of graph structure learning (GSL) (Zhu et al., 2022) is to make GNNs more robust against adversarial perturbations of the graph or to learn a graph for data where there is no graph to start with (Jin et al., 2020; Chen et al., 2020b; Wang et al., 2024; Zhu et al., 2024). Lv et al. (2023) build a neighborhood graph over features and interpolate between it and the input graph, which is a form of alignment. Unlike our method, they do not use spectral information, are unable to deal with noisy features and are only suitable for homophilic graphs where similarity-based connection rules apply. Even though both our work and GSL consider noisy graph settings, they are conceptually very different. We do not add noise to graph datasets which corresponds to a perturbation rate of 00 (“clean data”) in GSL nomenclature. Instead, we acknowledge that in every real world dataset, there is noise in the graph structure and the node features, and one manifestation of this noise is a misalignment of their leading eigenspaces. We then use this to rewire the graph (and denoise features) so as to improve the overall node classification performance of downstream GNNs. Naturally, GSL methods have difficulties to improve over baselines in this setting (Jin et al., 2020; Dong and Kluger, 2023). Our method, on the other hand, is not designed to handle strong perturbations and therefore cannot compete with GSL methods developed for specifically this purpose.
@@
Graph Regularization. Laplacian regularization (Ando and Zhang, 2006), originally stemming from semi-supervised representation learning, has been adapted by recent methods (Yang et al., 2021; Ma et al., 2021a) to also improve the performance of GNNs. An extra loss term is added during the GNNs training, which contains additional information about the graph structure to reduce oversmoothing. Their main limiting factor is the underlying homophiliy assumption: It is assumed that connected nodes are more likely to share the same label.
@@ 5 Conclusion and Limitations
Still, we should emphasize that preprocessing rewiring methods only use the graph, not the features. A direct comparison is thus not fair, as JDR cannot be used at all without features. Nonetheless, we have proposed the first generally applicable algorithm for cases where such information is available.
On a meta level, our results suggest that noise in real-world graphs is an important limiting factor for the performance of GNNs. It would therefore also be interesting to see whether feature-agnostic rewiring from a denoising perspective, for example using link prediction, could be used to improve the downstream performance. A related idea that we tried, but could not get to work well, is combining existing geometric rewiring algorithms with JDR. Intuitively, there should be a way to benefit from both removing noise and facilitating computation, but we have to leave that for future work.
We also note that most current rewiring methods can be applied to graph level tasks, while JDR is limited to node classification tasks, as it is derived via cSBM, which only considers community structures. It is therefore a fundamental research question how to extend the cSBM idea to graph-level problems. Given the already ubiquitous impact of cSBM on GNN research, this could potentially have a major impact on the field.
@@ A.1.2 Rotational invariance of Alignment
Here, we show that the alignment measure (3) is invariant to rotations of the subspaces for non-unique eigenvalues. Let $A\in\mathbb{R}^{N\times N}$ be the adjacency matrix and $X\in\mathbb{R}^{N\times F}$ the feature matrix. Let the eigendecomposition of $A$ be
@@ A.1.3 Computational Complexity
The complexity of JDR results mainly from the SVD and eigendecomposition, which is of order $\mathcal{O}(FN\text{min}(F,N))$ for SVD and $\mathcal{O}(N^{3})$ for the eigendecomposition ($F=N$). Since we only need the leading $k$ eigenvectors this reduces to $\mathcal{O}(FNk)$. If the matrix is additionally sparse as is often the case for real-world graphs with binary node features this reduces further to $\mathcal{O}(z(A)k)$ where $z(A)$ is the number of non-zero elements in $A$. If we assume that neither the average degree $d$ nor $k$ is scaled by $N$ (or $F$), the complexity becomes $\mathcal{O}(N)$.
@@ A.2 Contextual Stochastic Block Models
SBMs and GMMs are landmark theoretical models for studying clustering, classification problems and developing algorithmic tools. The cSBM (Deshpande et al., 2018), a combination of the two, has become a key model for studying node classification problems on graphs, inspiring numerous designs of GNNs like GPRGNN (Chien et al., 2021), GIANT (Chien et al., 2022) or ASGC (Chanpuriya and Musco, 2022). Many theoretical studies of node-level GNN problems are based on the cSBM, e.g. on double descent (Shi et al., 2024), neural collapse (Kothapalli et al., 2023), OOD generalization (Baranwal et al., 2021), or oversmoothing (Wu et al., 2023). Beyond being a standard synthetic benchmark, the cSBM is also used to to verify hypotheses about GNNs (Ma et al., 2022; Luan et al., 2023).

### v3

@@ 3 Experiments
Table 1: Comparison of state-of-the-art preprocessing rewiring approaches. Note that we refer to the computational complexity per iteration. $N$ denotes the number of nodes, $m$ the number of edges and $d_{\text{max}}$ is the maximum node degree. Additional details on the complexity of JDR are given in Appendix A.1.3; detailed runtime comparisons are in Appendix A.1.4
@@
. Method Add edge Remove edge Use Features Heterophilic? Complexity DIGL [Gasteiger et al., 2019] $\mathcal{O}(N)$ FoSR [Karhadkar et al., 2023] $\mathcal{O}(N^{2})$ BORF [Nguyen et al., 2023] $\mathcal{O}(md_{\text{max}}^{3})$ JDR (Ours) $\mathcal{O}(N)$
@@
We extensively evaluate JDR on both synthetic data generated from the cSBM and real-world benchmark datasets. We follow experimental setting from Chien et al. [2021] and evaluate JDR for semi-supervised node classification with different downstream GNNs. We also adopt their data splits, namely the sparse splitting $2.5\%/2.5\%/95\%$ for training, validation and testing, respectively, or the dense splitting $60\%/20\%/20\%$. For the general experiments, we perform $100$ runs with different random splits. For the scalability experiments, we use the experimental settings of the respective works [Lim et al., 2021, Platonov et al., 2023]. We report the average accuracy and the $95\%$-confidence interval calculated via bootstrapping with 1000 samples. All experiments are reproducible using the code provided.
@@
Baselines. Following recent works on rewiring, we use graph convolution network (GCN) [Kipf and Welling, 2017] as our downstream GNN. To obtain a more comprehensive picture, we additionally evaluate the performance on the more recent generalized PageRank graph neural network (GPRGNN) [Chien et al., 2021]. We compare our algorithm with the state-of-the-art rewiring methods first-order spectral rewiring (FoSR) [Karhadkar et al., 2023], batch Ollivier-Ricci flow (BORF) [Nguyen et al., 2023] and diffusion improves graph learning (DIGL) [Gasteiger et al., 2019]. FoSR approximates which edges should be added to maximize the spectral gap to reduce oversquashing. BORF adds edges in regions of negative curvature in the graph, which indicate bottlenecks that can lead to an oversquashing of the messages passed along these edges. A positive curvature indicates that there are so many edges in this area that messages could be oversmoothed, which is why edges are removed here. We compare computational and implementation aspects of JDR and baselines in Table 1. On the cSBM, we compare to an optimal baseline, namely the approximate message passing-belief propagation (AMP-BP) algorithm [Duranthon and Zdeborová, 2023]. AMP-BP is asymptotically optimal (in the large dimension limit) for unsupervised or semi-supervised community detection in the cSBM. It relies on knowing the distribution of the cSBM and is thus not applicable to real-world graphs with unknown characteristics and complex features.
@@
Hyperparameters. Unless stated otherwise, we use the hyperparameters from Chien et al. [2021] for the GNNs and optimize the hyperparameters of JDR using a mixture of grid and random search on the validation set. We use the top-$64$ values of $\tilde{\bm{A}}$ to enforce sparsity and interpolation to update the features. For DIGL, FoSR and BORF, we tune their hyperparameters using a grid search, closely following the given parameter range from the original papers. For all hyperparameter searches we use GCN and GPRGNN as the downstream models on $10$ runs with different random splits. A detailed list of all hyperparameters can be found in Appendix A.7 or in the code repository.
@@ 3.2 Results on Real-World Data
We evaluate JDR on five common homophilic benchmarks datasets, namely the citation graphs Cora, CiteSeer, PubMed [Sen et al., 2008] and the Amazon co-purchase graphs Computers and Photo [McAuley et al., 2015]. For heterophilic datasets, we rely on the Wikipedia graphs Chameleon and Squirrel [Rozemberczki et al., 2021], the WebKB datasets Texas and Cornell used in Pei et al. [2020] and the actor co-occurence network Actor [Tang et al., 2009]. To show the scalability of JDR on larger heterophilic datasets, we further report the results for the Yandex Q user network Questions [Platonov et al., 2023] and the social networks Penn94 and Twitch-Gamers [Lim et al., 2021]. Further details about all datasets are in Appendix A.4. Following Chien et al. [2021], we evaluate the homophilic datasets in the sparse splitting, staying close to the original setting of Kipf and Welling [2017] and the heterophilic datasets in dense splitting [Pei et al., 2020]. The remaining larger graphs are evaluated using their original splits. For further results and splits, see Appendix A.5.
@@
Heterophilic Datasets. The results in Table 3 show that GCN +JDR can catch up significantly compared to GPRGNN, but GPRGNN +JDR generally performs better. This is in line with the findings for the heterophilic cSBM. DIGL performs well on Actor, Texas and Cornell despite its inherent homophily assumption. The reason for this is the chosen smoothing kernel, which results in a graph that is evenly connected everywhere with small weights. GCN then largely ignores the graph and thus performs very similarly to an MLP, which performs already quite well on these datasets [Chien et al., 2021]. However, this fails for GPRGNN, which can make better use of the weak, complex graph structures. FoSR and BORF also improve performance here in most cases, but they are outperformed by JDR in all cases, often by a large margin. The out-of-memory error on Squirrel for BORF results from its computational complexity of $\mathcal{O}(md_{\text{max}}^{3})$, because the dataset has a large number of edges $m$ and a high maximum node degree $d_{\text{max}}$.
@@ 4 Relation to Prior Work
Graph Rewiring. Recent work show that even when a graph correctly encodes interactions it may not be an effective computational graph for a GNN due to conditions such as oversquashing [Alon and Yahav, 2021, Di Giovanni et al., 2023] and oversmoothing [Chen et al., 2020a]. Recently many methods have been proposed to address this, notably graph rewiring methods. They can be divided into preprocessing and end-to-end methods. Preprocessing methods rewire the graph using geometric and spectral properties, including curvature [Topping et al., 2022, Nguyen et al., 2023, Fesser and Weber, 2024, Bober et al., 2024], expansion [Deac et al., 2022, Banerjee et al., 2022], effective resistance [Black et al., 2023, Shen et al., 2024], and spectral gap [Karhadkar et al., 2023]. Conceptually related is diffusion-based rewiring [Gasteiger et al., 2019] that smooths the graph with a diffusion kernel. This can be interpreted as graph denoising, but is only suitable for homophilic graphs. Our approach is related to rewiring but with several key differences (see Table 1). Our rewiring strategy aims to denoise the graph (rather than control some geometric property) with the goal to improve downstream performance, while the classical rewiring literature focuses on optimizing the graph for message passing computations.
@@
Early end-to-end methods randomly drop edges during training to reduce oversmoothing [Rong et al., 2020]. Subsequent work [Gutteridge et al., 2023, Qian et al., 2024] incorporates latent features to dynamically rewire the graph. Ji et al. [2023] use the estimated labels of a GNN to rewire the graph during training of the same GNN. Giraldo et al. [2023] use curvature information for dynamic rewiring. Graph Transformers [Dwivedi and Bresson, 2021, Rampasek et al., 2022] aim to overcome oversquashing in GNNs via global attention. In order to handle large graphs, these works still need to revert to sparse, non-global attention [Gabrielsson et al., 2023, Shirzad et al., 2023]. Unlike preprocessing methods, end-to-end methods cannot output an improved graph which restricts their interpretability and reusability.
Graph Denoising. There is extensive literature on denoising signals on graphs using graph filters [Chen et al., 2014, Ma et al., 2021b, Liu et al., 2022]. However, we are interested in modifying the structure of the graph itself (rewiring), in a way that can benefit any downstream algorithm. Dong and Kluger [2023] recently proposed a new metric to measure graph noise that correlates well with GCN performance. Based on this, they develop a method for graph denoising via self-supervised learning and link prediction. We discuss the relation to our work in detail in Appendix A.5.7 and also evaluate our rewired graphs using their ESNR metric there. More broadly, link prediction [Zhang and Chen, 2018, Pan et al., 2022] can be seen as a tool for graph denoising; this perspective has been applied, for instance, to denoising neighborhood graphs arising in molecular imaging [Debarnot et al., 2022].
@@
Graph Structure Learning. The aim of graph structure learning (GSL) [Zhu et al., 2022] is to make GNNs more robust against adversarial perturbations of the graph or to learn a graph for data where there is no graph to start with [Jin et al., 2020, Chen et al., 2020b, Wang et al., 2024, Zhu et al., 2024]. Lv et al. [2023] build a neighborhood graph over features and interpolate between it and the input graph, which is a form of alignment. Unlike our method, they do not use spectral information, are unable to deal with noisy features and are only suitable for homophilic graphs where similarity-based connection rules apply. Even though both our work and GSL consider noisy graph settings, they are conceptually very different. We do not add noise to graph datasets which corresponds to a perturbation rate of 00 (“clean data”) in GSL nomenclature. Instead, we acknowledge that in every real world dataset, there is noise in the graph structure and the node features, and one manifestation of this noise is a misalignment of their leading eigenspaces. We then use this to rewire the graph (and denoise features) so as to improve the overall node classification performance of downstream GNNs. Naturally, GSL methods have difficulties to improve over baselines in this setting [Jin et al., 2020, Dong and Kluger, 2023]. Our method, on the other hand, is not designed to handle strong perturbations and therefore cannot compete with GSL methods developed for specifically this purpose.
@@
Graph Regularization. Laplacian regularization [Ando and Zhang, 2006], originally stemming from semi-supervised representation learning, has been adapted by recent methods [Yang et al., 2021, Ma et al., 2021a] to also improve the performance of GNNs. An extra loss term is added during the GNNs training, which contains additional information about the graph structure to reduce oversmoothing. Their main limiting factor is the underlying homophiliy assumption: It is assumed that connected nodes are more likely to share the same label.
@@ 5 Conclusion and Limitations
The main limitation of JDR is that it cannot be used without node features. The preprocessing rewiring methods that we compare with do not have this limitation as they only use the graph for rewiring, but in turn, they cannot take advantage of features. Since JDR is the first method to jointly denoise the graph and the features, there are no other methods to which it could be directly compared. Our experiments thus highlight what advantage features bring to rewiring.
Furthermore, our results suggest that noise in real-world graphs is an important limiting factor for the performance of GNNs. It would be interesting to see whether feature-agnostic rewiring from a denoising perspective, for example using link prediction, could be used to improve the downstream performance. A related idea that we tested but could not get to work well is to combine existing geometric rewiring algorithms with JDR. Intuitively, there should be a way to benefit from both removing noise and facilitating computation, but we have to leave that exploration for future work.
We also note that most current rewiring methods can be applied to graph level tasks, while JDR is currently limited to node classification. It is an open question how to extend the cSBM idea to graph-level problems.
Acknowledgments
JL, CS and ID were supported by the European Research Council (ERC) Starting Grant 852821—SWING.
@@ A.1.2 Rotational invariance of Alignment
We show that the alignment measure (3) is invariant to rotations of the subspaces for non-unique eigenvalues. Let $A\in\mathbb{R}^{N\times N}$ be the adjacency matrix and $X\in\mathbb{R}^{N\times F}$ the feature matrix. Let the eigendecomposition of $A$ be
@@ A.1.3 Computational Complexity
The complexity of JDR results mainly from the SVD and eigendecomposition, which is of order $\mathcal{O}(FN\text{min}(F,N))$ for SVD and $\mathcal{O}(N^{3})$ for the eigendecomposition ($F=N$). Since we only need the leading $k$ eigenvectors this reduces to $\mathcal{O}(FNk)$. If the matrix is additionally sparse as is often the case for real-world graphs with binary node features this reduces further to $\mathcal{O}(\text{nnz}(A)k)$ where $\text{nnz}(A)$ is the number of non-zero elements in $A$. Since usually neither the average degree $d$ nor $k$ is scaled by $N$, the complexity actually scales with $\mathcal{O}(N)$.
A.1.4 Runtime Comparison
In addition to computational complexity, we report measurements of running time of the different algorithms. We run JDR and baseline methods on the real-world datasets, using GCN as downstream model. All algorithms are run on Nvidia A100 with $80$GB and we time their Python processes. We emphasize that we did not explicitly optimize the timing code and we kept the outputs and logging turned on. But this influences all methods in the same way so the relative comparisons are meaningful. The results in Table 5 do not show a clear “winner”. The ambiguity is especially visible on the large heterophilic graphs, where JDR is slower than DIGL, but significantly faster than FoSR on two datasets. On the Twitch-gamers dataset, on the other hand, it is faster than DIGL but not as fast as FoSR. The main reason for this is that different hyperparameters choices of the rewiring methods lead to dramatically different run times (even when applying the same method on the same dataset). For example on Computers, JDR is very fast since it only requires $3$ denoising iterations, compared to $15$ on Citeseer. The same holds true for FoSR which only require $5$ iterations on Twitch-gamers, but $700$ on Questions. So if one wants to optimize for speed, one should constrain the hyperparameters of the methods that significantly impact execution speed. Of course, there is a trade-off between any such constraint and accuracy, as our experiments on the denoising iterations of JDR in Figure 16 (b) and Figure 17 (b) in Appendix A.6 show.
Table 5: Timing experiments in seconds for different rewiring methods using GCN as downstream GNN. Smaller is better. We record the time of the preprocessing and training and evaluating the GNN on $100$ random splits. The results do not show a clear winner; JDR generally requires a comparable or less time compared to baselines. For more discussion see A.1.4.
Dataset Base DIGL FoSR BORF JDR
Cora 182 228 187 201 246
Citeseer 258 360 291 290 433
PubMed 291 692 416 897 858
Computers 274 444 516 718 465
Photo 213 299 220 801 330
Chameleon 372 239 396 483 545
Squirrel 1263 302 1282 - 1659
Actor 166 286 171 225 319
Texas 163 200 169 174 203
Cornell 167 202 164 240 208
Questions 93 804 11053 - 3707
Penn94 164 232 3198 - 1779
Twitch-gamers 1579 6729 1618 - 5165
@@ A.2 Contextual Stochastic Block Models
SBMs and GMMs are landmark theoretical models for studying clustering, classification problems and developing algorithmic tools. The cSBM [Deshpande et al., 2018], a combination of the two, has become a key model for studying node classification problems on graphs, inspiring numerous designs of GNNs like GPRGNN [Chien et al., 2021], GIANT [Chien et al., 2022] or ASGC [Chanpuriya and Musco, 2022]. Many theoretical studies of node-level GNN problems are based on the cSBM, e.g. on double descent [Shi et al., 2024], neural collapse [Kothapalli et al., 2023], OOD generalization [Baranwal et al., 2021], or oversmoothing [Wu et al., 2023]. Beyond being a standard synthetic benchmark, the cSBM is also used to to verify hypotheses about GNNs [Ma et al., 2022, Luan et al., 2023].